# AI Knows Best? The Paradox of Expertise, AI-Reliance, and Performance in Educational Tutoring Decision-Making Tasks

### v1

@@ 4.2 Time Spent Comparison
Highlight + Reasoning 11.86 (7.12) 12.10 (6.04) -0.140 0.889 -0.037
The results are presented in Figure 3 and Table 4. We conducted a repeated measure 2-way ANOVA to examine the main effects of Inline Highlighting and Text Reasoning, as well as their interaction with the data after IQR filtering. The analysis revealed a significant main effect respectively for Inline Highlighting ($F(1,57)=25.74,p<.001,\eta_{p}^{2}=.311$) and Text Reasoning ($F(1,57)=19.12,p<.001,\eta_{p}^{2}=.251$). Additionally, there was a significant interaction effect between Highlighting and Reasoning ($F(1,57)=18.26,p=.001,\eta_{p}^{2}=.243$). This interaction suggests that the combined effect of Inline Highlighting and Text Reasoning does not lead to a substantial increase in time; rather, their effects partially offset each other. Moreover, we found no significant difference in time spent between participants from Tutor and Prolific.
4.3 Over-reliance across groups
The results comparing over-reliance percentages for Tutors and Prolific participants across different task conditions are presented in Figure 4 and Table 5. We conducted a one-way ANOVA to investigate differences in over-reliance rates across conditions, revealing a highly significant effect $(F(5,570)=16.4074,p<.0001)$. Post-hoc analysis using Tukey’s HSD showed that, for questions where the AI was incorrect, the control group without any AI labels or suggestions exhibited significantly lower over-reliance rates than participants who received AI suggestions (all $p_{\text{adj}}<.05$). These findings indicate that the presence of wrong AI suggestions led participants to make wrong decisions, whether in the form of labels, Text Reasoning, or Inline Highlighting.
Figure 4: Average Over-Reliance Percentage, where users were wrong when AI’s suggestion was wrong, across different conditions and groups. It can be observed that participants from Prolific are more prone to over-reliance. Error bar is Standard Error.
Table 5: Over-Reliance T-test Comparison Results Between Tutor and Prolific Participants.
@@ 4.3 Over-reliance across groups
Highlighting 0.654 (0.279) 0.792 (0.215) -2.696 0.0083**
Highlighting + Reasoning 0.718 (0.253) 0.839 (0.175) -2.702 0.0082**
@@
When further comparing over-reliance between different groups of users, as shown in Table 5 and Figure 4, we found that under the conditions of Highlighting, Highlighting + Reasoning, AI Label Only, and for overall average, Prolific participants significantly exhibited more over-reliance compared to the Tutor. However, under the conditions without any AI suggestions and with only Reasoning explanations, the two groups had no significant difference in over-reliance.
4.4 Under-reliance across groups
The results comparing under-reliance percentages for Tutors and Prolific participants across different conditions are presented in Figure 5 and Table 6.
Table 6: Under-Reliance Comparison T-test Results Between Tutor and Prolific Participants.
@@ 4.4 Under-reliance across groups
Highlighting 0.202 (0.155) 0.099 (0.111) 3.742 0.0003**
Highlighting + Reasoning 0.160 (0.161) 0.083 (0.112) 2.721 0.0078**
@@
Figure 5: Average Under-Reliance Percentage, where users were wrong even if AI’s suggestions are correct, across different Conditions and Groups. It can be observed that tutors are more prone to under-reliance. The error bar is Standard Error.
A one-way ANOVA examined differences in under-reliance percentages across conditions and revealed a highly significant effect ($F(5,570)=7.9017,p<.001$). Subsequent Tukey’s HSD post-hoc tests indicated that, for questions that AI was correct, the control group without labels exhibited significantly higher error (under-reliance) rates compared to conditions with AI suggestions ($p_{\text{adj}}=.05$). This finding suggests that participants tend to trust the AI, and thus, when the AI is correct, they tend to respond correctly, too.
@@
When further comparing under-reliance between different groups of users, we found that under the conditions of Inline Highlighting, Highlighting + Reasoning, and AI Label Only, the Tutor group significantly exhibited more under-reliance compared to the Prolific group. However, under the conditions without any AI suggestions and with Text Reasoning explanations, the two groups had no significant difference in under-reliance.
5 Discussion
5.0.1 Human-AI Collaboration vs. “AI-Only” Approaches.
First, while participants do benefit from AI-generated explanations—particularly when both text reasoning and highlighting are provided—the improvement remains modest. Across most conditions, overall human-AI team accuracy lags behind the AI-only baseline 0.8. This result suggests that, even when collaborating with AI in judging the correctness of tutor responses, humans fail to achieve a complementary effect in decision-making tasks. That is, unlike some earlier studies [14, 13, 2], we did not observe a complementary effect whereby experts collaborating with AI outperformed either the AI by itself or the human expert working alone.
@@ 5 Discussion
From a practical standpoint, these results pose a question about the necessity of human intervention in educational decision-making tasks. It may be the case that an ensemble of multiple AIs (e.g., a majority vote from different Large Language Models) surpasses both humans alone and human-AI teams. This trend raises concerns about efficiency and accountability: if an AI ensemble makes the final decision better than a human, and humans will be biased by AI, who is held accountable for errors? Future work should thus explore reliable, “responsible” automation methods that can ensure accountability—whether through transparency measures, oversight processes, or clear guidelines for error remediation.
5.0.2 Reliance on AI: Novice Over-Reliance vs. Expert Under-Reliance
@@
Our findings are consistent with previous research showing that, despite its potential advantages, human judgment often leads to errors through either under-reliance or over-reliance on AI [2, 19, 16]. Moreover, we found that in the context of evaluating tutor responses, text reasoning exerts the more pronounced influence on both sides of the trust dynamic. Specifically, textual reasoning reduces under-reliance (users become more convinced when AI is correct) but increases over-reliance (users are less likely to reject incorrect AI suggestions). In contrast, inline highlighting on its own did not show a significant effect. In contrast, inline highlighting alone did not show a significant effect. These findings aligned with previous research as well that inline highlighting can reduce reliance [2, 16, 14, 13].
Yet, neither explanation strategy significantly improved performance nor reached a complementary effect, which differs from previous research [2, 19, 14, 13]. This suggests that current approaches still have room for improvement. Future research in AI-assisted educational decision-making should thus explore more adaptive explanation styles and strategies that mitigate both over- and under-reliance while still leveraging the benefits of AI assistance.
@@
Moreover, we observed notable differences between experienced tutors (“experts”) and non-tutors recruited from Prolific (“novices”). Novices were more inclined to trust the AI’s suggestions, leading to higher overall accuracy but greater over-reliance when the AI was wrong. Conversely, they demonstrated fewer under-reliance errors when the AI was correct. In other words, novices consistently “followed” the AI—an approach that performs well overall if the AI maintains a high baseline accuracy. Experts, by contrast, occasionally under-relied on correct AI suggestions, reflecting skepticism or a sense of confidence in their own expertise. These findings align with prior research indicating that novices tend to rely on AI more [14]. Future efforts in designing human-AI collaboration systems should, therefore, account for users’ varying levels of expertise and adapt explanation strategies to balance trust and skepticism accordingly.
5.0.3 Limitations and Generalizability
@@
Moreover, our task involved a single type of decision-making task (evaluating the appropriateness of tutor praise). Other educational tasks—like open-ended response evaluation or essay grading—may pose different challenges for AI explanation and user reliance. Our results, therefore, highlight the need for broader investigations into how AI-assisted decision-making scales across diverse educational contexts. Identifying consistent patterns in over-reliance and under-reliance can guide better interface design, explanation style, and training protocols, leading to improved collaboration between humans and AI in education.
6 Conclusion
Our findings reinforce a complex reality: although AI explanations can boost performance under certain conditions—particularly for novices—designers must carefully navigate the interplay between over-reliance, under-reliance, and practical constraints such as time. In addition, the style of explanations must be deliberately selected and tailored. To fully harness AI’s potential, future research should investigate more on adaptive explanation approaches in AI-assisted decision-making within educational settings, explore ensemble methods that enhance performance, and establish nuanced accountability frameworks to mitigate potential reliance.

### v2

@@ 4.1. Decision-Making Performance
Highlight + Text Reasoning 11.86 (7.12) 12.10 (6.04) -0.140 0.889 -0.037
Figure 3. Average Time Spent by Condition and Group. The data is obtained after excluding outliers using the 1.5 interquartile range (IQR) method. Participants spend more time on conditions that have AI explanations. The error bar is Standard Error.
When further comparing performance between the two groups across conditions, as shown in Table 1, we found that under the conditions of Inline Highlighting, Highlighting + Reasoning, and for overall average, the Prolific group significantly outperformed the Tutor group. The Prolific group also marginally significantly outperformed the Tutor group in the AI Label Only case. However, under the conditions without any AI suggestions and with only Text Reasoning explanations, the two groups had no significant difference.
4.2. Time Spent comparison
The results are presented in Figure 3 and Table 2. We conducted a repeated-measure 2-way ANOVA to examine the main effects of Inline Highlighting and Text Reasoning, as well as their interaction with the data after IQR filtering. The analysis revealed a significant main effect respectively for Inline Highlighting ($F(1,57)=25.74,p<.001,\eta_{p}^{2}=.311$) and Text Reasoning ($F(1,57)=19.12,p<.001,\eta_{p}^{2}=.251$). Additionally, there was a significant interaction effect between Highlighting and Reasoning ($F(1,57)=18.26,p=.001,\eta_{p}^{2}=.243$). This interaction suggests that the combined effect of Inline Highlighting and Text Reasoning does not lead to a substantial increase in time; rather, their effects partially offset each other. Moreover, we found no significant difference in time spent between participants from Tutor and Prolific.
4.3. Over-reliance
Table 3. Over-Reliance T-test Comparison Results Between Tutor and Prolific Participants.
@@ 4.3. Over-reliance
Inline Highlighting 0.654 (0.279) 0.792 (0.215) -2.696 0.0083**
Inline Highlighting + Text Reasoning 0.718 (0.253) 0.839 (0.175) -2.702 0.0082**
@@
Figure 4. Average over reliance percentage (users were wrong when the AI was wrong) across conditions and groups. Participants from Prolific show higher over reliance. Error bars represent standard error.
The results comparing over-reliance percentages for Tutors and Prolific participants across different task conditions are presented in Figure 4 and Table 3. We conducted a one-way ANOVA to investigate differences in over-reliance rates across conditions, revealing a highly significant effect $(F(5,570)=16.4074,p<.0001)$. Post-hoc analysis using Tukey’s HSD showed that, for questions where the AI was giving incorrect suggestions, participants who received AI suggestions showed significantly more errors compared to the baseline condition (all $p_{\text{adj}}<.05$). These findings indicate that the presence of wrong AI suggestions led participants to make wrong decisions, whether in the form of labels, Text Reasoning, or Inline Highlighting.
@@
When further comparing performance between the two groups across conditions, as shown in Table 3 and Figure 4, we found that under the conditions of Highlighting, Highlighting + Reasoning, AI Label Only, and for overall average, Prolific participants significantly exhibited more over-reliance compared to the Tutor. However, under baseline and Text Reasoning conditions, the two groups showed no significant difference.
4.4. Under-reliance
Table 4. Under-Reliance Comparison T-test Results Between Tutor and Prolific Participants.
@@ 4.4. Under-reliance
Inline Highlighting 0.202 (0.155) 0.099 (0.111) 3.742 0.0003**
Inline Highlighting + Text Reasoning 0.160 (0.161) 0.083 (0.112) 2.721 0.0078**
@@
Figure 5. Average Under-Reliance Percentage, where users were wrong even if AI’s suggestions are correct, across different Conditions and Groups. It can be observed that tutors are more prone to under-reliance. The error bar is Standard Error.
The results comparing under-reliance percentages for Tutors and Prolific participants across different conditions are presented in Table 4 and Figure 5. A one-way ANOVA revealed a significant effect of condition on under-reliance ($F(5,570)=7.9017,p<.001$). Subsequent Tukey’s HSD post-hoc tests indicated that, for questions for which AI was correct, participants who received AI suggestions had significantly less error rate compared to the baseline condition (all $p_{\text{adj}}=.05$). This finding suggests that participants tend to trust the AI, and thus, when the AI is correct, they tend to respond correctly.
@@
When further comparing performance between the two groups across conditions, we found that under the conditions of Inline Highlighting, Highlighting + Reasoning, and AI Label Only, the Tutor group significantly exhibited more under-reliance compared to the Prolific group. However, under baseline and Text Reasoning conditions, the two groups had no significant difference.
5. Discussion
5.1. Human-AI Collaboration vs. “AI-Only” Approaches.
@@ 5. Discussion
First, while participants do benefit from AI-generated explanations, particularly when both text reasoning and highlighting are provided, the improvement remains modest. Across most conditions, overall human-AI team accuracy still falls short of the AI-only baseline of 0.8. This suggests that even when collaborating with AI to judge the correctness of tutor responses, humans fail to achieve a complementary effect in decision-making tasks. In other words, unlike some earlier studies (morrison2023impact; morrison2023evaluating; bansal2021does), we did not observe a complementary effect in which experts working with AI outperformed either the AI alone or the human expert alone. From a practical standpoint, these results pose a question about the necessity of human intervention in educational decision-making tasks. It may be the case that an ensemble of multiple AIs (e.g., a majority vote from different Large Language Models) surpasses both humans alone and human-AI teams. This trend raises concerns about efficiency and accountability: if an AI makes the final decision better than a human, and humans will be biased by AI, who is held accountable for errors? Future work should therefore explore reliable and “responsible” automation methods that ensure accountability, whether through transparency measures, oversight processes, or clear guidelines for error remediation.
5.2. Text Reasoning, Highlighting, and Reliance Patterns Difference in Expertise
@@
Our findings are consistent with previous research showing that, despite its potential advantages, human judgment often leads to errors through either under-reliance or over-reliance on AI (bansal2021does; vereschak2021evaluate; si2023large). Moreover, we found that in the context of evaluating tutor responses, text reasoning exerts the more pronounced influence on both sides of the trust dynamic. Specifically, textual reasoning reduces under-reliance (users become more convinced when AI is correct) but increases over-reliance (users are less likely to reject incorrect AI suggestions). In contrast, inline highlighting on its own did not show a significant effect. These findings align with previous research, which also shows that inline highlighting can reduce reliance (bansal2021does; si2023large; morrison2023impact; morrison2023evaluating). However, neither explanation strategy significantly improved performance nor reached a complementary effect, which differs from previous research (bansal2021does; vereschak2021evaluate; morrison2023impact; morrison2023evaluating). This suggests that current approaches still have room for improvement. Future research in AI-assisted educational decision-making should thus explore more adaptive explanation styles and strategies that mitigate both over- and under-reliance while still leveraging the benefits of AI assistance.
Furthermore, our results indicate that text reasoning has a significantly stronger influence on reliance than inline highlighting. Text reasoning significantly reduced under-reliance but also increased over-reliance, showing that explanatory content can meaningfully shift user judgments and must therefore be designed with particular care. In contrast, inline highlighting had less measurable effect, suggesting a need for future work on more refined visual cues that guide attention without increasing bias.
@@
We also observed notable differences between non-tutors recruited from Prolific (“novices”) and experienced tutors (“experts”). Novices showed a stronger tendency to trust AI suggestions, which increased their overall accuracy but also made them more vulnerable to over-reliance when the AI was incorrect. This pattern aligns with prior findings (morrison2023impact). Experts, by contrast, were more skeptical and sometimes overrode correct AI suggestions, leading to under-reliance errors. This pattern suggests that in educational decision-making contexts, domain expertise may increase confidence in one’s own judgment, even in cases where the AI is reliably correct and does better than a human alone. This pattern differs from earlier studies in non-educational contexts (morrison2023impact). Future efforts in designing human-AI collaboration systems in education should, therefore, account for users’ varying levels of expertise and self-efficacy, and adapt explanation strategies to balance trust and skepticism accordingly.
5.3. Implications
In addition to performance measures, indicators such as over-reliance, under-reliance, and time cost can serve as process-level indicators for modeling human-AI collaboration. These traces can be logged and analyzed to understand how explanation designs calibrate trust and efficiency across user groups. This framing positions human–AI decision-making as analyzable evidence for learning analytics, motivating future work on adaptive dashboards and monitoring tools for AI-assisted decision support. Additionally, these indicators can also potentially be used to help train humans to become better decision-makers in AI-assisted contexts.
5.4. Limitations and Future Works
@@
Moreover, our study focused on a single type of decision-making task (evaluating the appropriateness of tutor praise). Other educational tasks, such as open-ended response evaluation or essay grading, may present different challenges for AI explanation and user reliance. Our findings, therefore, underscore the need for broader investigations into how AI-assisted decision-making applies across diverse educational contexts. Identifying consistent patterns of over-reliance and under-reliance can inform better interface design, explanation styles, and training protocols, ultimately fostering more effective collaboration between humans and AI in education.
6. Conclusion
Our findings highlight a complex reality: although AI explanations can improve performance under certain conditions, particularly for novices, designers must carefully manage the balance between over-reliance, under-reliance, and practical constraints such as time. By framing reliance and efficiency measures as learning analytics constructs, our study shows how AI-assisted decision-making can generate analyzable process data that informs both theory and practice. To fully harness AI’s potential in education, future research should explore adaptive explanation strategies grounded in learning analytics, investigate how these indicators can be embedded into real-time decision-support dashboards, and establish accountability frameworks that integrate human expertise with data-driven insights. In this way, human-AI collaboration can move beyond interface design to contribute directly to the broader learning analytics cycle of data collection, analysis, interpretation, and informed action.
Acknowledgements.
This research was supported by the Generative AI + Education Tools R&D Seed Grant at Carnegie Mellon University. We also thanks to the National Science Foundation (award CNS-2213791) for partial support of this work.